At Euclid High School, the number of students taking the AMC10 was 60 in
2002, 66 in 2003, 70 in 2004, 76 in 2005, and 78 in 2006, and is 85 in 2007.
Between what two consecutive years was there the largest percentage increase?
(A) 2002 and 2003 (B) 2003 and 2004 (C) 2004 and 2005 (D) 2005 and 2006
(E) 2006 and 2007

Answer by Theo(13342)

60 = 2002
66 = 2003
70 = 2004
76 = 2005
78 = 2006
85 = 2007

rate of increase for each year would be the value in that year minus the value in the previous year divided by the value in the previous year.

rate of increase in 2003 over 2002 = 6/60 = .1
rate of increase in 2004 over 2003 = 4/66 = .06
rate of increase in 2005 over 2004 = 6/70 = .085
rate of increase in 2006 over 2005 = 2/76 = .026
rate of increase in 2007 over 2006 = 7/78 = .089

percent equals 1005 * rate.

biggest increase percent is 2003 over 2002 = 10%